OpenStudy (anonymous):

1-1/4(2-(3-x)/(5)) ≤ (8-x)/(3) I will give a medal.

OpenStudy (debbieg):

Is it this? \[\Large 1-\dfrac{ 1 }{ 4 }\left( 2-\dfrac{\left( 3-x \right)}{5} \right)\le \dfrac{ 8-x }{ 3 }\]